OpenStudy (anonymous):

Add. 4 2/9 and 9 1/9

OpenStudy (anonymous):

40/3

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it 13 3/9

OpenStudy (anonymous):

first u add 4 +9 then u add 2 +1 and not the denominators

OpenStudy (anonymous):

13 3/9 = 13 1/3 = 40/3